Find the general solutions of the following differential equations using D-operator methods:

2.1 (D2 - 2D + 5)y = x+5

2.2 (D+4)2 x = sinh4t


1
Expert's answer
2022-04-06T17:18:34-0400

"2.1:\\\\\\left( D^2-2D+5 \\right) y=x+5\\\\\\phi \\left( D \\right) =D^2-2D+5=\\left( D-1 \\right) ^2+2^2\\\\\\left( \\left( D-1 \\right) ^2+2^2 \\right) y=x+5\\\\The\\,\\,solution\\,\\,of\\,\\,the\\,\\,homogeneous\\,\\,equation\\,\\,is\\\\y=Ae^x\\cos 2x+Be^x\\sin 2x\\\\Particular\\,\\,solution:\\\\y=C_0+C_1x:\\\\\\left( D^2-2D+5 \\right) \\left( C_0+C_1x \\right) =x+5\\\\-2C_1+5\\left( C_0+C_1x \\right) =x+5\\\\\\left\\{ \\begin{array}{c}\t-2C_1+5C_0=5\\\\\t5C_1=1\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tC_1=0.2\\\\\tC_0=1.5\\\\\\end{array} \\right. \\\\y=1.5+0.2x+Ae^x\\cos 2x+Be^x\\sin 2x\\\\2.2:\\\\\\left( D+4 \\right) ^2x=\\sinh 4t\\\\e^{-4t}D^2e^{4t}x=\\sinh 4t\\\\D^2e^{4t}x=e^{4t}\\sinh 4t\\\\D^2e^{4t}x=\\frac{1}{2}e^{8t}-\\frac{1}{2}\\\\De^{4t}x=\\int{\\left( \\frac{1}{2}e^{8t}-\\frac{1}{2} \\right) dt}=\\frac{1}{16}e^{8t}-\\frac{1}{2}t+C_1\\\\e^{4t}x=\\int{\\left( \\frac{1}{16}e^{8t}-\\frac{1}{2}t+C_1 \\right) dt}=\\frac{1}{128}e^{8t}-\\frac{1}{2t^2}+C_1t+C_2\\\\x=\\frac{1}{128}e^{4t}-\\frac{1}{2t^2}e^{-4t}+C_1te^{-4t}+C_2e^{-4t}"